OpenStudy (anonymous):

decide whether the composite functions, fog and gof, are equal to x. f(X)= x+4/3, g(X)= 3x-4

OpenStudy (anonymous):

f o g = (3x - 4) + 4 all over 3 which would be 3x / 3 = x g o f = 3(x+4/3) - 4 x + 4 - 4 = x Yes they are